Status Report

NASA ISS On-Orbit Status 3 September 2008

By SpaceRef Editor
September 3, 2008
Filed under , , ,
NASA ISS On-Orbit Status 3 September 2008

All ISS systems continue to function nominally, except those noted previously or below.

Before breakfast, the CDR, FE-1 and FE-2 completed another periodic session of the Russian biomedical routine assessment PZEh-MO-7/Calf Volume Measurement (seventh for CDR & FE-1, fifth for FE-2). [Calf measurements (left leg only) are taken with the IZOG device, a custom-sewn fabric cuff that fits over the calf, using the knee and lower foot as fixed reference pints, to provide a rough index of deconditioning in zero-G and effectiveness of countermeasures.]

Volkov & Kononenko each spent ~1.5 hrs on the TVIS for the periodic Russian PZE-MO-3 test for physical fitness evaluation, their second time, using the TVIS in unmotorized (manual control) mode and wearing the Kardiokassette KK-2000 belt with three chest electrodes. [The fitness test, controlled from the RSE-Med laptop, yields ECG (electrocardiogram) readings to the KK-2000 data storage device, later downlinked via the Regul (BSR-TM) payload telemetry channel. Before the run, the KK-2000 was synchronized with the computer date/time readings. For the ECG, the crewmembers worked out on the treadmill, first walking 3 min. up to 3.5 km/h, then running at a medium pace of 6.5 km/h, followed by the maximum pace not exceeding 10 km/h, then walking again at gradually decreasing pace.]

Volkov also performed the periodic maintenance of the active Russian BMP (Harmful Impurities Removal System) by starting the "bake-out" cycle to vacuum on absorbent bed #1 of the regenerable dual-channel filtration system. The process will be terminated at ~5:15pm EDT and Bed #2 regeneration performed tomorrow. [Regeneration of each of the two cartridges takes about 12 hours and is conducted only during crew awake periods. The BMPs regeneration cycle, normally done every 20 days, is currently performed four times more frequently to remove any lingering Freon-218 from the cabin atmosphere (last time: 8/28 & 29)];

Chamitoff worked on the outfitting of the U.S. Regenerative ECLSS, to install the new oxygen (O2) port on the PD4 standoff in the Lab for the WRS (Water Recovery System), after venting the new CHeCS (Crew Health Care Systems) rack O2 supply hose to prevent contamination, then relocating the umbilical to the new port, purging it and the CHeCS O2 port and configuring the setup for an overnight leak check. This activity was deferred on 8/25 when the FE-2 could not locate, for a while, the missing NOPA (N2 O2 Purge Adapter) hardware required for purging to evacuate any air in the system. [The CHeCS rack will be relocated in the near future in the U.S. Lab from the D4 position to the S4 position in order to install the WRS-1 rack during STS-126/ULF2. Since there will be no O2 connection to the CHeCS rack at the S4 location, the new O2 port must be installed, also to interface with the Regenerative ECLSS system to maintain a functional O2 port in the D4 bay of the Lab.]

After filling the CHeCS, WRS-1 and WRS-2 MTL (Moderate Temperature Loop) supply jumpers with ITCS (Internal Thermal Control System) coolant at two Lab positions, set up yesterday, the FE-2 today disconnected and removed the remaining MTL jumper. [The hoses need to be filled prior to mating them to their respective racks. The WRS will be used to recycle wastewater into potable water. The Regenerative ECLSS will be required for the future six-person occupancy of the station.]

Kononenko and Volkov conducted an audit of free stowage space in the RS (Russian Segment), i.e., FGB, SM (Service Module) and DC1 (Docking Compartment), addressing specific questions uplinked in a questionnaire on 8/29.

The CDR also retrieved the three copies of the 1J Warning Book from Lab, SM and FGB, replaced two pages (on Loop A/B thermal pump outlet temperature limits) with new material, then restowed the books. The old pages were discarded as trash.

The FE-1 loaded the Russian laptops RSK1, RSK2, RSE-med, RSS1 and RSE1 with new antivirus software from DVD, scanned the hard disks, taking documentary Nikon photos of the resulting message displays, and consulted with ground specialists via S-band tagup.

The FE-2 initiated the ghost loading of the SAMS (Space Acceleration Measurement System) hard drive with new image structures, to be terminated about 7 hrs later before bedtime.

Oleg continued the current round of periodic preventive maintenance of cabin ventilation systems in the RS, today cleaning the Group A fan grilles in the SM.

Gregory had 2:10 hrs reserved for the first part of CGSE (Common Gas Supply Equipment) troubleshooting in the JAXA JPM (JEM Pressurized Module). [Working on the SSV (Self-Shutoff Valve) in the CGSE Valve Unit, Chamitoffs task today was to recover the SSV, i.e., restore functionality, then isolate the SAIBO rack from the system CO2 line by disconnecting the QD (Quick Disconnect) on the UIP (Utility Interface Panel/Z-Panel) ands repressurize the system CO2 line, to be monitored by SSIPC (Space Station Integration & Promotion Center) at Tsukuba/Japan for two days to determine if either the system or the payload is the CO2 leak source. On 8/8, the CBEF (Cell Biology Experiment Facility) in the SAIBO rack exhibited a sudden unexpected pressure drop at the CGSE CO2 high-pressure sensor, from 3000 kPa to 500 kPa within 1.5 minutes. Specialists believe that the SSV closed itself unexpectedly, making it a potential contributor to the sudden pressure decrease. Also suspected is leakage from the CO2 supply line. Part 2 of troubleshooting, on 9/5, will depend on the result of the overnight leak check.]

Later today (~3:30pm EDT), Oleg Kononenko will support a 20-min. systems test of the newly installed ATV PCE (Automated Transfer Vehicle/Proximity Communications Equipment, Russian: MBRL), specifically the BUAP Antenna Feeder Unit and the PU control console for the ATV, preparatory to Fridays undocking of Jules Verne.

Sergey Volkov conducted the periodic audit/inspection of available photo/video equipment in the RS, in particular four camcorders (including three SONY HVR-Z1J digital high-definition video cameras), one DVCAM, adapter brackets and mounting plates.

Greg Chamitoff had ~1:50 hr set aside for continuing to unpack and stow cargo delivered on STS-124/1J/Discovery on 6/2.

Sergey took care of the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM. [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]

The CDR also completed the daily IMS (Inventory Management System) maintenance, updating/editing its standard delta file including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).

On the Russian Matryoshka-R (RBO-3-2) radiation payload in the DC1, Oleg checked out a memory card and activated the AST Spectrometer for data takes. [After retrieving a used PCMCIA (Portable Computer Memory Card International Adapter) card, ALC-951, from stowage, Oleg activated the AST twice to verify the cards labels and orientation, then re-inserted ALC-951 in the spectrometer and activated the AST again.]

The FE-1 took the periodic photographs of the plants growing in the BIO-5 Rasteniya-2 ("Plants-2") Lada-13 greenhouse, using the Nikon D2 photo camera with F=17-55 mm lens, then downlinked the images via Regul BSR-TM. The regular status check of the payload which researches growth and development of plants (barley) under spaceflight conditions was conducted as per suggestion by the discretionary time permitting task list.

In the U.S. Lab, Sergey Volkov performed a controlled shut-down of the EHS VOA (Environmental Health System-Volatile Organic Analyzer), with the ground power-cycling its RPC (Remote Power Controller).

Gregory completed another one of the periodic inspections of the RED (Resistive Exercise Device) canister cords and accessories, currently done every two weeks.

The station residents conducted their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2), TVIS treadmill (CDR/MO-3, FE-1/MO-3), RED resistive exercise device (FE-1, FE-2) and VELO bike with bungee cord load trainer (FE-1). Later, Kononenko transferred the exercise data file to the MEC (Medical Equipment Computer) laptop for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

As generally every day now, starting at ~9:00am and running until 3:00pm, the US CDRA (Carbon Dioxide Removal Assembly) was activated intermittently for two half-cycles to control ppCO2 levels. This configuration for the daily ops does not require connecting & disconnecting the ITCS (Internal Thermal Control System) cooling loop. [A forward plan is in work for cycling the CSV (CO2 Selector Valve) to prevent its sticking.]

ISS Wireless LAN Transition: Yesterday, after beginning the transition of the JSL (Joint Station LAN) Network to the new WAPs (Wireless Access Points) in Node-1 & Node-2 and loading the PDAs (Personal Data Assistants) with new applications for wireless (RF) comm, hardware and software configuration issues prevented Greg Chamitoff from completing the scheduled survey of wireless LAN (Local Area Network) connectivity aboard, intended to assess wireless LAN for Node-3 network interfaces (vs. wired interfaces). The survey will have to be re-scheduled.

JAXA Kibo GLA Failures: A total of six GLAs (General Luminaire Assemblies) have now been reported failed in the JPM. There are still ten GLAs functioning, and there is no impact for crew activity. In the JLP (JEM Logistics Pressurized Segment), two GLAs have failed, with only one GLA remaining. Removal of the failed light fixtures and replacement with spares is being planned for both Kibo modules.

CEO (Crew Earth Observations) photo targets uplinked for today were S. Mozambique, Africa (ISS had a nadir pass over this site, and overlapping mapping frames taken along track were requested. The region is undergoing rapid development, and imagery of the existing state of land cover will be useful to track changes to regional ecosystems over time), Hurricane Hanna, Atlantic Ocean (Dynamic Event. This tropical cyclone was predicted to regain hurricane strength by the time of this pass [10:47am EDT]. Looking ahead and to the right of track for cloud banding and eye features. The storm will be passing over the Bahamas), High Central Andean Glaciers, S. America (weather was predicted to be mostly clear over the eastern slopes of the Andes. Looking to the right of track for glaciers and icefields on the upper mountain slope – imagery of these glaciers is useful for tracking changes in ice extent and volume. Overlapping mapping frames were requested as ISS paralleled the Andes from NW to SE), and Beni River dynamics, Bolivia (the Beni River rises in the eastern Andes range, and flows for almost 1600 km [994 miles] across Bolivia. General context imagery of the river was requested to guide later higher resolution imagery of the river channel and banks. The regional of interest for photography is east of the Andes. Greg was to look for Lake Titicaca as a landmark feature).

CEO photography can be studied at this Gateway website: (as of 3/1/08, this database contained 757,605 views of the Earth from space, with 314,000 from the ISS alone).

Week 20/21 Scheduled Main Activities:

  • Thu. (9/4): SOTR-KOB1 maintenance; SRVK-BRPK maintenance; BMP ch.2 regen; CWC audit; SLEEP init.; PCS s/w load-ghosting; IWIS reprogram; ATV: activate/remove equipment & SSVP docking system/close hatches/take photo+video; TVS-MPEG Ku-band test.
  • Fri. (9/5): ATV Undock (5:27pm EDT); PCS reboot; microbial sampling; CGSE troubleshoot.
  • Sat. (9/6): FFQ; IWIS deact.
  • Sun. (9/7): Station cleaning; DCP bypass cable install; PFC (FE-2).
  • Mon. (9/8): SLEEP; Acoustic Survey; BMP regen ch.1; POC-DOUG revw; SSRMS/ESP-3; Hatch seal inspect; CGSE troubleshoot; EMER OBT; ATV MBRL/PCE uninstall; Iridium phone recharge.
  • Tue. (9/9): SLEEP; Acoustic Survey; BMP regen ch.2; TORU OBT; O-OHA; EMU batt. maint.; BCAT-3 fam.; PMCs; JAXA AQUA-Ink Ball exp.

ISS Orbit (as of this morning, 7:36am EDT [= epoch]):
Mean altitude — 353.5 km
Apogee height — 358.0 km
Perigee height — 348.9 km
Period — 91.61 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.0006736
Solar Beta Angle — 9.6 deg (magnitude increasing)
Orbits per 24-hr. day — 15.72
Mean altitude loss in the last 24 hours — 44 m
Revolutions since FGB/Zarya launch (Nov. 98) — 56085

Significant Events Ahead (all dates Eastern Time, some changes possible.):
09/05/08 — ATV1 undocking, from SM aft port (~5:27pm); independent flight
09/08/08 — Progress M-64/29P de-orbit (~4:45pm)
09/10/08 — Progress M-65/30P launch (~3:49:45pm)
09/12/08 — Progress M-65/30P docking (SM aft, ~5:08pm DM)
09/29/08 — ATV de-orbit (nighttime re-entry for observation)
10/01/08 — NASA 50 Years (official)
10/08/08 — STS-125/Atlantis Hubble Space Telescope Service Mission 4 (SM4)
10/11/08 — Progress M-65/30P undocking (from SM aft)
10/12/08 — Soyuz TMA-13/17S launch (~3:03am EDT; Lonchakov, Fincke, Garriott)
10/14/08 — Soyuz TMA-13/17S docking (FGB nadir port, ~4:51am)
10/23/08 — Soyuz TMA-12/16S undocking (DC1 nadir) or 10/24?
11/10/08 — STS-126/Endeavour/ULF2 launch MPLM Leonardo, LMC
11/12/08 — STS-126/Endeavour/ULF2 docking
11/20/08 — ISS 10 Years
11/25/08 — Progress M-65/30P undocking & deorbit
11/26/08 — Progress M-66/31P launch
11/30/08 — Progress M-66/31P docking
02/09/09 — Progress M-66/31P undocking & deorbit
02/10/09 — Progress M-67/32P launch
02/12/09 — Progress M-67/32P docking
02/12/09 — STS-119/Discovery/15A launch S6 truss segment
02/14/09 — STS-119/Discovery/15A docking
02/24/09 — STS-119/Discovery/15A undocking
02/26/09 — STS-119/Discovery/15A landing (nominal)
03/25/09 — Soyuz TMA-14/18S launch
03/27/09 Soyuz TMA-14/18S docking (DC1)
04/05/09 — Soyuz TMA-13/17S undocking
04/07/09 — Progress M-67/32P undocking & deorbit
05/15/09 — STS-127/Endeavour/2J/A launch – JEM EF, ELM-ES, ICC-VLD
05/25/09 — Soyuz TMA-15/19S launch
05/27/09 — Six-person crew on ISS (following Soyuz 19S docking)
07/30/09 — STS-128/Atlantis/17A MPLM(P), last crew rotation
10/15/09 — STS-129/Discovery/ULF3 – ELC1, ELC2
12/10/09 — STS-130/Endeavour/20A Node-3 + Cupola
02/11/10 — STS-131/Atlantis/19A MPLM(P)
04/08/10 — STS-132/Discovery/ULF4 ICC-VLD, MRM1
05/31/10 — STS-133/Endeavour/ULF5 ELC3, ELC4 (contingency).

SpaceRef staff editor.